OpenStudy (anonymous):

A customer went to a garden shop and bought some potting soil for $17.50 and 4 shrubs. The total bill was $53.50. Write and solve an equation to find the price of each shrub

OpenStudy (anonymous):

\[\frac{1}{4} (53.5\, -17.5)=9 \]

OpenStudy (anonymous):

What you need to do in these questions is take out the essential data as follows: potting soil = $17.50 number of shrubs = 4 total bill = $53.50. let x = a shrub therefore the total amount he spent was $17.50 + 4(x) = $53.50 Now we just use the rules of algebra to find the price of a shrub: First, subtract $17.50 from both sides as follows: $17.50 (-$17.50) + 4(x) = $ 53.50 - ($17.50) 4(x) = $36.00 dividing across by 4 we get 4(x)/4 = $36.00/4 x = $9